OpenStudy (anonymous):

1/8 times 3/4 = ? 3/12 4/24 3/32 3/24

OpenStudy (anonymous):

To multiply fractions you just multiply straight across, SO: 1x3/8x4 the just simplify.

OpenStudy (anonymous):

*then

OpenStudy (ganpat):

1/8 * 3/ 4 = 3/32

OpenStudy (anonymous):

1/8*3/4= 3/32 so option 3 is the answer